OpenStudy (anonymous):

Solve m plus one-fourth equals three-fourths. m equals negative one-half m equals one-fourth m equals one-half m equals one

OpenStudy (anonymous):

m+1/4=3/4

OpenStudy (anonymous):

Come on @i_need_help123 you can do this.

OpenStudy (anonymous):

"Solve m plus one-fourth equals three-fourths" becomes "m plus" means m + "one-fourth" means 1/4 "equals three-fourths" means = 3/4 put it all together and you get m + 1/4 = 3/4

OpenStudy (anonymous):

srry not really paying attention im doing something as im doing this

OpenStudy (anonymous):

Well pay attention for a second so we can help you. What do you think the answer is with the information we gave you?

OpenStudy (anonymous):

@i_need_help123 this websites main purpose (in my opinion) Is to help understanding and improve your grades. Not give out answer's. So @i_need_help123 it would help if you pay attention and stop expecting somebody to give you the answer. So please what do you think the answer is so I can check/review and help you. ( on both your questions I'm in today)

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

what do you think the answer is?